OpenStudy (anonymous):

A number minus 20 times its reciprocal is 1. Find the number.

OpenStudy (anonymous):

x-20/x=1

OpenStudy (kira_yamato):

Undefined for all real values of x. Because when you solve the equation, you get 0 = 20 or 0 = -20, and you end up with nonsense

OpenStudy (s):

n - 20 * 1/n = 1 n - 20/n = 1 times everything by n n^2 - 20 = n n^2 - n - 20 = 0 (n-5)*(n+4) = 0 n1 = 5; n2 = -4 But if the question is: (n - 20) * 1/n = 1 then (n-20)/n = 1 times everything by n n-20 = n n - n = 20 0 = 20 which is nonsense O_o because 0 is not equal to 20 O_o